OpenStudy (anonymous):

solve the triangle A=46(degrees) ,a=31 ,b=27 Please help, i keep getting the wrong answers

OpenStudy (anonymous):

IDK I got 2 but that's

OpenStudy (anonymous):

the options are B = 38.8°, C = 115.2°, c ≈ 34.3 B = 38.8°, C = 95.2°, c ≈ 25.7 B = 38.8°, C = 95.2°, c ≈ 42.9 Cannot be solved

OpenStudy (anonymous):

I'll try it again sorry

OpenStudy (anonymous):

no problem thanks

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

did you get an answer yet?

OpenStudy (anonymous):

c=42.91

OpenStudy (anonymous):

cosA=(b^2+c^2 - a^2)/(1*bc)

OpenStudy (anonymous):

yeah i tried the equation and i would get like 0.75. But thanks :)

OpenStudy (anonymous):

/2bc not /1bc hope it helps

OpenStudy (anonymous):

dunno i used that equation but my calculator gave me 42.91729 my pleasure

OpenStudy (anonymous):

* A = 46 degrees If it's a right triangle, use a^2+b^2=c^2, or a trig function. If it's obtuse, use a trig function, or law of sines/cosines 2 possible triangles could be formed. Basically imagine side a intersecting side c at a different point. In either triangle Angle A is still 46° and side a = 31 and side b = 27.

OpenStudy (anonymous):

IDK, I hope this stuff helps

OpenStudy (anonymous):

sorry I tried